Find the exact value of the trigonometric expression given that
sin(u) = −
3
5
,
where
3π/2 < u < 2π,
and
cos(v) =
15
17
,
where
0 < v < π/2.
cot(u + v)

Answer by Alan3354(69443)

I think you mean:
Find the exact value of the trigonometric expression given that
sin(u) = −3/5 where 3π/2 < u < 2π,
and cos(v) = 15/17 where 0 < v < π/2.
--------
cot(u + v)
========================
cot = cos/sin
--------
Find the cos and the sine of the sum.